Let B={u1= (1, 1, -1) u2= (1, 0, 1) u3 = (3, 2, 0)} be a basis for R3. Find the dual basis of B


1
Expert's answer
2021-01-27T01:08:11-0500

The linear function may be expressed in the form

"{\phi}_1(x,y,z)=a_1x+a_2y+a_3z" , "\phi_2(x,y,z)=b_1x+b_2y+b_3z ,"

"\phi_3(x,y,z)=c_1x+c_2y+c_3z"


By the definition of dual basis ,

"\phi_i(v_j)=0 \ for \ i\neq j" but "\phi_i(v_j)=1" for "i=j" .

We find "\phi_1" by setting "\phi_1(u_1)=1, \phi_1(u_2)=0,\phi_1(u_3)=0" . This yields

"\phi_1(u_1)=a_1+a_2-a_3=1,\phi_1(u_2)=a_1+a_3=0 \ , \ \phi_1(u_3)=3a_1+2a_2=0"

Solving the system of equation yields

"a_1=2 \ ,\ a_2=-3 \ , \ a_3=-2"

Thus , "\phi_1(x,y,z)=2x_1-3x_2-2x_3"

We find "\phi_2" by setting , "\phi_2(u_1)=0 \ , \ \phi_2(u_2)=1 \ , \ \phi_3(u_3)=0" , This yields

"\phi_1(u_1)=b_1+b_2-b_3=0,\phi_1(u_2)=b_1+b_3=1 \ , \ \phi_1(u_3)=3b_1+2b_2=0"

Solving the system of equation yields

"b_1=2 \ , \ b_2=-3 \ , \ b_3=-1"

Thus , "\phi_2(x,y,z)=2x - 3x_2 - x_3"

Finally , we find "\phi_3" by setting "\phi_1(u_1)=c_1+c_2-c_3=0,\phi_1(u_2)=c_1+c_3=0 \ , \ \phi_1(u_3)=3c_1+2c_2=1"

Solving the system of equation yields , "c_1=0 \ ,\ c_2=\frac{1}{2} \ , \ c_3=0" .

Thus , "\phi_3(x,y,z)=\frac{1}{2}x_2" .